OpenStudy (anonymous):

3b divided by c ;b = 12 ,c =4

OpenStudy (anonymous):

3b/c so 3*12/4 = 36/4=9

OpenStudy (anonymous):

i dont under stand that -_-

OpenStudy (anonymous):

Just plug in the different numbers to the variables...

OpenStudy (anonymous):

the answer is 9. you takve 3 multiplied by 12 to get 36 then dvided 36 by 4 and get the answer 9.